OpenStudy (anonymous):

Consider triangle ABC with vertices A(4,8), B(0,0) and C(6,0). AQ and BP are two altitudes of triangle ABC. AQ and BP meet at H. (a) Find the equations of AQ and BP. (b) Find the coordinates of H. (c) If CR is the third altitude of ABC, show that CR also passes through H.

hartnn (hartnn):

could you draw a corresponding figure ?

OpenStudy (anonymous):



hartnn (hartnn):

to get the equation of AQ, u need its slope, to get its slope u need slope of BC, which is ?

hartnn (hartnn):

can u find slope of BC ?

hartnn (hartnn):

or, did u realize that AQ is a vertical line ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

slope of BC=6 ?

hartnn (hartnn):

no, slope of BC = (0-0)/(6-0) = 0 right ? u know slope formula ?

OpenStudy (anonymous):

oh I forgot that the line which is horizontal and the slope is zero. ok...

hartnn (hartnn):

right, and vertical line's equation is of the form ?

OpenStudy (anonymous):

vertical line.....slope is undefined right?

hartnn (hartnn):

yup,. the equation of vertical line is of the form x=c c is the x-co-ordinate.

OpenStudy (anonymous):

got it.

hartnn (hartnn):

got the equation of both the lines ?

OpenStudy (anonymous):

AQ=4 ---> x=4 BP=1/4 y=(1/4)x+0 y=(1/4)x

hartnn (hartnn):

yup, now u know how to find co-ordinates of H ?

OpenStudy (anonymous):

wait a minute, let me write that down...

OpenStudy (anonymous):

H(4,1)

hartnn (hartnn):

yup.

OpenStudy (anonymous):

got it

hartnn (hartnn):

can u find equation of CR in same way ?

OpenStudy (anonymous):

 let me think.........

hartnn (hartnn):

find slope of AB first.....

OpenStudy (anonymous):

finding.... mAB=8-0/4-0=2

OpenStudy (anonymous):

slope of CR * slope of AB=-1 mCR= -1 / 2

hartnn (hartnn):

yep, now equation of CR

OpenStudy (anonymous):

I cannot get the equation...............

hartnn (hartnn):

y-y1 = m(x-x1) m =-1/2 x1,y1 = 6,0

OpenStudy (anonymous):

oh I need to substitute the coordinates of H into the formula, right?

hartnn (hartnn):

not H, put co-ordinates of C

OpenStudy (anonymous):

got it.

hartnn (hartnn):

because u need to prove that H lies on CR, so u can't use H

OpenStudy (anonymous):

oh yes, I find it just now....thank you :)

hartnn (hartnn):

good, welcome ^_^